Two point electric charges q2 = 5q1 and q1 are at a distance of 3 meters and apply a repulsive force of 0.02 N to each other, and how many microclones?

1
Expert's answer
2022-03-13T18:47:04-0400

The Coulomb's law says

"F=k\frac{q_1q_2}{r^2}=k\frac{5q_1^2}{r^2}"

Hence

"q_1=\sqrt{Fr^2/(5k)}\\
=\sqrt{0.02*3^2/(5*9*10^9)}=2*10^{-6}\:\rm C=2\:\mu C"

"q_2=5q_1=5*2\:\mu C=10\:\mu C"